# Singularity of the axisymmetric stagnation-point-like solution within a cylinder of the 3D Euler incompressible fluid equations

### v1

@@ 3.2 Initial Stretching Rate Continuous at the Minimum Point
To gain deeper insight into the relationship between the initial stretching rate and the singular behavior of the ASCII fluid, we consider a more general case in which the initial stretching rate $\gamma_{0}(r_{0})$ is assumed to be continuous in the neighborhood of its minimum point. For simplicity, we also assume that $\gamma_{0}(r_{0})$ has only one minimum point—hence, only one potential blowup point.
@@
This definition ensures that $n_{\pm}$ characterise the leading-order algebraic rate at which $\gamma_{0}(r_{0})$ approaches its minimum from either side. In particular, $n_{\pm}<\infty$ corresponds to power-law behavior near $r_{-}$, while $n_{\pm}=\infty$ indicates that $\gamma_{0}(r_{0})$ approaches f{f} faster than any power law—such as in the case of Gevrey-class regularity at $r_{-}$.
@@
Notice that, because of the formula (34), the absence of a finite-time singularity (namely, the case $T_{*}=\infty$) necessarily requires $\lim_{t\to T_{*}}\Lambda_{1}(t)=\infty$. Therefore, a careful analysis of the convergence or divergence properties of $\Lambda_{1}(t)$ and $\Lambda_{2}(t)$ as $t\to T_{*}$ is essential. These integrals become improper as $S(t)\to S_{*}=-1/{f}$, due to the vanishing denominator at the minimum point $r_{0}=r_{-}$. In the following subsections, we examine their behavior in detail under various assumptions on $_{-}, n+n_{+}$, and $r_{-}$. We note that the presence of the max function in the definition of the coefficient $n$ in equation (58) is directly related to the fact that the original system’s regularity ($T_{*}=\infty$) requires a divergent $\Lambda_{1}$, which is realised in terms of the maximum between $n_{-}$ and $n_{+}$, as we will see below in equation (65).
@@ 3.2.1 Convergence and Divergence of the Integrals $\Lambda_{1}(t)$ and $\Lambda_{2}(t)$
$\quad+\frac{{f}r_{-}}{[\lambda_{-}-\varepsilon_{-}(\zeta_{0})]^{2}}\lim_{\nu\to 0}\int_{\nu}^{r_{-}}r_{0}^{-2n_{-}}\mathrm{d}r_{0}+\frac{{f}r_{-}}{[\lambda_{+}-\varepsilon_{+}(\zeta_{2})]^{2}}\lim_{\nu\to 0}\int_{\nu}^{R-r_{-}}r_{0}^{-2n_{+}}\mathrm{d}r_{0}$
$\quad-\frac{{f}}{[\lambda_{-}-\varepsilon_{-}(\zeta_{1})]^{2}}\lim_{\nu\to 0}\int_{\nu}^{r_{-}}r_{0}^{1-2n_{-}}\mathrm{d}r_{0}+\frac{{f}}{[\lambda_{+}-\varepsilon_{+}(\zeta_{3})]^{2}}\lim_{\nu\to 0}\int_{\nu}^{R-r_{-}}r_{0}^{1-2n_{+}}\mathrm{d}r_{0},$
@@
The convergence or divergence of $\Lambda_{1}(t)$ and $\Lambda_{2}(t)$ is determined by the values of $r_{-}$, $n_{-}$, and $n_{+}$. The following tables 1 and 2 summarises respectively the conditions under which the integral $\Lambda_{1}(t)$ and $\Lambda_{2}(t)$ converges or diverges:
$n\in$ $(0,\frac{1}{2})$ $\frac{1}{2},1) [$1,2)[1,2) $2,+\infty)$
$r_{-}=0$ Converges Converges Converges Diverges
$r_{-}\neq 0$ Converges Converges Diverges Diverges
Table 1: Convergence behavior of $\Lambda_{1}(t)$ as $t\to T_{*}$ for different values of $n$ and $r_{-}$.
$n\in$ $(0,\frac{1}{2})$ $\frac{1}{2},1) [$1,2)[1,2) $2,+\infty)$
$r_{-}=0$ Converges Converges Diverges Diverges
$r_{-}\neq 0$ Converges Diverges Diverges Diverges
Table 2: Convergence behavior of $\Lambda_{2}(t)$ as $t\to T_{*}$ for different values of $n$ and $r_{-}$.
This classification reveals a clear dependence on the exponents $n_{-}$ and $n_{+}$, as well as on the location of the minimum point $r_{-}$. In particular, in terms of the coefficient $n$ defined in equation (58):
@@ 3.2.2 Asymptotic Behavior of the Integrals $\Lambda_{1}(t)$ and $\Lambda_{2}(t)$: Divergent Case
where $p\geq 0$ and $q\geq 1$ are integers. Using this definition, the integrals $\Lambda_{1}(t)$ and $\Lambda_{2}(t)$ can be rewritten in terms of these components as follows:
@@ 3.2.3 Asymptotic Behavior of the Integrals $\Lambda_{1}(t)$ and $\Lambda_{2}(t)$: Convergent Case
Based on these results, we summarise the asymptotic behaviors of $\Lambda_{1}(t)$ and $\Lambda_{2}(t)$ in tables 3 and 4, where the constants $G_{n}$ and $L_{j}$ are defined as:
@@
The left portion of the table corresponds to the case $r_{-}=0$, while the right portion applies to $r_{-}\in(0,R]$. These results, together with those derived in equations (74) and (75) for the special cases where $n_{-}n_{+}=\infty$, allow us to determine the structure of singular solutions in the ASCII fluid model.
@@
From this analysis, we see that, regarding the asymptotic behaviour of $\Lambda_{1}$ and $\Lambda_{2}$ as $t\to T_{*}$:
  • a)
    For small values of $n_{\pm}$, the integrals remain close to their limiting values $\Lambda_{1*}$ and $\Lambda_{2*}$, indicating convergent behaviour;
  • b)
    As $n$ increases, divergent behaviour appears and becomes more pronounced;
  • c)
@@
These asymptotic results will serve as the foundation for determining the regularity or blowup nature of the vorticity coordinate $(\gamma,\omega)$ and the vertical pathline $z(z_{0},r_{0},t)$ in the subsequent analysis. Recall that a convergent behaviour of $\Lambda_{1}$ as $t\to T_{*}$ implies a finite-time singularity of the solution to the original fluid equations (i.e. $T_{*}<\infty$).
@@ 3.2.4 Asymptotic Behaviors of the Vorticity Coordinate
    Central minimum: when $n$ is finite and the minimum point lies at the centre: $r_{-}=0$;
@@
    Non-central minimum: when $n$ is finite and the minimum point lies away from the centre: $r_{-}\in(0,R]$.
@@
These quantities appear frequently in the asymptotic expressions for $\omega$, $\gamma$, and $z$, and their definitions encapsulate the dependence on the cylinder radius $R$, the location of the blowup point $r_{-}$, and the exponents $n_{-},n_{+}$ that characterise the local behavior of $\gamma_{0}(r_{0})$ near its minimum.
The subsequent analysis will focus on how the vorticity coordinate and vertical pathline behave near the singularity time under these different scenarios. In particular, we examine the scaling laws of the blowup and determine whether the divergence occurs in finite time or not.
Firstly, table 5 summarises the case of flat minimum, where $n_{-}n_{+}=+\infty$, meaning that at least one side (either left or right) of the blowup point remains flat within a neighborhood. The result showing an infinite singularity time $T_{*}=\infty$ indicates that the ASCII fluid remains regular for all finite times. Consequently, it is not meaningful to discuss the asymptotic behavior of the flow near the singularity in this scenario.
@@
Secondly, we consider the case of central minimum, where $r_{-}=0$ and $n_{+}>0$ , corresponding to a flow that becomes singular at the axis of symmetry with the stretching rate exhibiting a local minimum at the minimum point $r_{0}=r_{-}$. These asymptotic behaviors are summarised in table 6. A critical threshold emerges at $n_{+}=4$, which we refer to as the blowup threshold. This means that the 3D Euler fluid under Lagrangian description remains regular if the local power-law approximation of the initial stretching rate satisfies $n_{+}\geq 4$, but develops a finite-time singularity when $n_{+}<4$.
@@
  • 2.
    Central Minimum: Asymptotic Behavior of Plane Vorticity $\omega$ and Vertical Pathline $z$
@@
    • iii)
      Moreover, the dependence on the cylinder radius is given by $\omega,z\propto R^{\frac{8}{4-n}-2}$, showing that larger radii accelerate the convergence of these quantities to zero.
@@
Thirdly, we consider the case of non-central minimum, where the minimum point of the initial stretching rate lies away from the axis of symmetry, i.e., $r_{-}\in(0,R]$, and $\gamma_{0}(r_{0})$ is continuous near this point with positive exponents $n_{\pm}\in\mathbb{R}_{+}$. The asymptotic behavior of the vorticity coordinate $(\omega,\gamma)$ and vertical pathline component $z$ in this scenario is summarised in table 7.
@@
These findings confirm that the local structure of the initial stretching rate $\gamma_{0}(r_{0})$ near its minimum plays a dominant role in determining whether and how the ASCII fluid develops a finite-time singularity. The critical thresholds $n=1$ (blowup-rate threshold) and $n=2$ (blowup threshold) define distinct regimes of behavior:
  • a)
@@
Now, we collect the following key observations regarding the relative strength of blowups, regardless of whether the blowup is central or non-central:
  • a)
@@
  • c)
    For $n\in[2,4)$, blowup occurs only when the minimum point lies at the centre; otherwise, for a non-central minimum, the solution remains regular;
  • d)
@@
regardless of whether the blowup occurs at the centre or away from it.
In addition to $r_{-}$, another parameter influencing the asymptotic behavior is $\delta$}, defined in equation (72), which quantifies the asymmetry between the left and right exponents $n_{-}$ and $n_{+}$ of $\gamma_{0}(r_{0})$ near its minimum. Since $\delta$} also appears in the denominator of $K_{1^{\pm}}$, it affects the blowup dynamics in the interval $n\in(1,2)$. Specifically, flows with asymmetric initial data—where $n_{-}\neq n_{+}$—exhibit faster blowup than those with symmetric profiles ($n_{-}=n_{+}$). In other words, an unsmooth or asymmetric profile of the initial stretching rate at the minimum point accelerates the formation of the singularity.
We emphasise a crucial physical distinction between central and non-central minimum scenarios. While the central minimum occurs on a point, the non-central minimum occurs on a ring, namely a degenerate case when compared to the point case. The ring case, due to the axisymmetry, has a perfectly “flat” profile in the azimuthal direction, which therefore changes the asymptotics in favour of more regular behaviour, hence the required parameter $n$ for regular behaviour is smaller compared with the central minimum case.
Finally we briefly address the question: What happens when the location of the minimum of the initial stretching rate is shared by several rings, or by a combination of rings and the central point? Namely, when the set $\{x\in\Omega\quad|\quad\gamma_{0}(x)=\inf\gamma_{0}\}$ is the union of two or more elementary sets of the form $\{x\in\Omega\quad|\quad|x|=r_{-}\}$, with $r_{-}\geq 0$. In such a case, the dominant ring (or centre point) is the one for which the integral $\Lambda_{1}$ grows faster asymptotically as $t\to T_{*}$. Thus, for example, if the minima of $\gamma_{0}$ were at a ring with $n=2$ and at the centre with $n=2$, then the asymptotic behaviour of the system would be dominated by the ring with $n=2$ because the integral $\Lambda_{1}$ would diverge due to the stretching rate’s profile at the ring, and thus the singularity time would still be $T_{*}=\infty$. As another example, a ring with $n=1$ combined with a centre point with $n=4$ would lead to no finite-time singularity because the integral $\Lambda_{1}$ would diverge due to the stretching rate’s profile at the centre.

### v2

@@ 3.2 Initial Stretching Rate Continuous at the Minimum Point
To gain deeper insight into the relationship between the initial stretching rate and the singular behavior of the cylindrical Gibbon flow, we consider a more general case in which the initial stretching rate $\gamma_{0}(r_{0})$ is assumed to be continuous in the neighborhood of its minimum point. For simplicity, we also assume that $\gamma_{0}(r_{0})$ has only one minimum point—hence, only one potential blowup point.
@@
This definition ensures that $n_{\pm}$ characterize the leading-order algebraic rate at which $\gamma_{0}(r_{0})$ approaches its minimum from either side. In particular, $n_{\pm}<\infty$ corresponds to power-law behavior near $r_{-}$, while $n_{\pm}=\infty$ indicates that $\gamma_{0}(r_{0})$ approaches f{f} faster than any power law—such as in the case of Gevrey-class regularity at $r_{-}$, (see, e.g., Rodino (1993),Kukavica and Vicol (2011)).
@@
Notice that, because of the formula (35), the absence of a finite-time singularity (namely, the case $T_{*}=\infty$) necessarily requires $\lim_{t\to T_{*}}\Lambda_{1}(t)=\infty$. Therefore, a careful analysis of the convergence or divergence properties of $\Lambda_{1}(t)$ and $\Lambda_{2}(t)$ as $t\to T_{*}$ is essential. These integrals become improper as $S(t)\to S_{*}=-1/{f}$, due to the vanishing denominator at the minimum point $r_{0}=r_{-}$. In the following subsections, we examine their behavior in detail under various assumptions on $_{-}, n+n_{+}$, and $r_{-}$. We note that the presence of the max function in the definition of the coefficient $n$ in equation (60) is directly related to the fact that the original system’s regularity ($T_{*}=\infty$) requires a divergent $\Lambda_{1}$, which is realized in terms of the maximum between $n_{-}$ and $n_{+}$, as we will see below in equation (67).
@@ 3.2.1 Convergence and Divergence of the Integrals $\Lambda_{1}(t)$ and $\Lambda_{2}(t)$
$\qquad+\frac{{f}r_{-}}{[\lambda_{-}-\varepsilon_{-}(\zeta_{0})]^{2}}\lim_{\nu\to 0}\int_{\nu}^{r_{-}}r_{0}^{-2n_{-}}\mathrm{d}r_{0}+\frac{{f}r_{-}}{[\lambda_{+}-\varepsilon_{+}(\zeta_{2})]^{2}}\lim_{\nu\to 0}\int_{\nu}^{R-r_{-}}r_{0}^{-2n_{+}}\mathrm{d}r_{0}$
$\qquad-\frac{{f}}{[\lambda_{-}-\varepsilon_{-}(\zeta_{1})]^{2}}\lim_{\nu\to 0}\int_{\nu}^{r_{-}}r_{0}^{1-2n_{-}}\mathrm{d}r_{0}+\frac{{f}}{[\lambda_{+}-\varepsilon_{+}(\zeta_{3})]^{2}}\lim_{\nu\to 0}\int_{\nu}^{R-r_{-}}r_{0}^{1-2n_{+}}\mathrm{d}r_{0},$
@@
The convergence or divergence of $\Lambda_{1}(t)$ and $\Lambda_{2}(t)$ is determined by the values of $r_{-}$, $n_{-}$, and $n_{+}$. Based on the limit as $t$ goes to $T_{*}$ of $\Lambda_{1}(t)$ and $\Lambda_{2}(t)$ in (66) and (67) under appropriate choices for $r_{-}$, $n_{-}$, and $n_{+}$, and using the fact that for any $\epsilon>0,\lim_{\nu\to 0}\int_{\nu}^{\epsilon}r_{0}^{\alpha}\mathrm{d}r_{0}=\infty$ when $\alpha\leq-1$, tables 1 and 2 summarize respectively the conditions under which the integral $\Lambda_{1}(t)$ and $\Lambda_{2}(t)$ converges or diverges. More precisely, this classification reveals a clear dependence on the exponents $n_{-}$ and $n_{+}$, as well as on the location of the minimum point $r_{-}$. In particular, in terms of the coefficient $n$ defined in equation (60):
@@
$n\in$ $(0,\frac{1}{2})$ $\frac{1}{2},1) [$1,2)[1,2) $2,+\infty)$
$r_{-}=0$ Converges Converges Converges Diverges
$r_{-}\neq 0$ Converges Converges Diverges Diverges
Table 1: Convergence behavior of $\Lambda_{1}(t)$ as $t\to T_{*}$ for different values of $n$ and $r_{-}$ , where $n$ is defined in (60).
$n\in$ $(0,\frac{1}{2})$ $\frac{1}{2},1) [$1,2)[1,2) $2,+\infty)$
$r_{-}=0$ Converges Converges Diverges Diverges
$r_{-}\neq 0$ Converges Diverges Diverges Diverges
Table 2: Convergence behavior of $\Lambda_{2}(t)$ as $t\to T_{*}$ for different values of $n$ and $r_{-}$ , where $n$ is defined in (60).
@@ 3.2.2 Asymptotic Behavior of the Integrals $\Lambda_{1}(t)$ and $\Lambda_{2}(t)$: Divergent Case
where $p\geq 0$ and $q\geq 1$ are integers. Using this definition, the integrals $\Lambda_{1}(t)$ and $\Lambda_{2}(t)$ expressed in (64) and (65) can be rewritten in terms of these components as follows:
@@ 3.2.3 Asymptotic Behavior of the Integrals $\Lambda_{1}(t)$ and $\Lambda_{2}(t)$: Convergent Case
Based on these results, we summarize the asymptotic behaviors of $\Lambda_{1}(t)$ and $\Lambda_{2}(t)$ in tables 3 and 4, where the constants $G_{n}$ and $L_{j}$ are defined as:
@@
The left portion of the table corresponds to the case $r_{-}=0$, while the right portion applies to $r_{-}\in(0,R]$. These results, together with those derived in equations (76) and (77) for the special cases where $n_{-}n_{+}=\infty$, allow us to determine the structure of singular solutions in the cylindrical Gibbon flow model.
@@
From this analysis, we see that, regarding the asymptotic behavior of $\Lambda_{1}$ and $\Lambda_{2}$ as $t\to T_{*}$:
  • For small values of $n_{\pm}$, the integrals remain close to their limiting values $\Lambda_{1*}$ and $\Lambda_{2*}$, indicating convergent behavior;
  • As $n$ increases, divergent behavior appears and becomes more pronounced;
@@
These asymptotic results will serve as the foundation for determining the regularity or blowup nature of the vorticity coordinate $(\gamma,\omega)$ and the vertical pathline $z(z_{0},r_{0},t)$ in the subsequent analysis. Recall that a convergent behavior of $\Lambda_{1}$ as $t\to T_{*}$ implies a finite-time singularity of the solution to the original fluid equations (i.e. $T_{*}<\infty$).
@@ 3.2.4 Asymptotic Behaviors of the Vorticity Coordinate
    Central minimum: when $n$ is finite and the minimum point lies at the center: $r_{-}=0$;
@@
    Non-central minimum: when $n$ is finite and the minimum point lies away from the center: $r_{-}\in(0,R]$.
@@
These quantities appear frequently in the asymptotic expressions for $\omega$, $\gamma$, and $z$, and their definitions encapsulate the dependence on the cylinder radius $R$, the location of the blowup point $r_{-}$, and the exponents $n_{-},n_{+}$ that characterize the local behavior of $\gamma_{0}(r_{0})$ near its minimum.
The subsequent analysis will focus on how the vorticity coordinate and vertical pathline behave near the singularity time under these different scenarios. In particular, we examine the blowup scaling laws and determine whether the divergence occurs in finite time or not.
a) Flat minimum. Firstly, table 5 summarizes the case of flat minimum, where $n_{-}n_{+}=+\infty$, meaning that at least one side (either left or right) of the blowup point remains flat within a neighborhood. The result showing an infinite singularity time $T_{*}=\infty$ indicates that the cylindrical Gibbon flow remains regular for all finite times. Consequently, it is not meaningful to discuss the asymptotic behavior of the flow near the singularity in this scenario.
@@
b) Central minimum. Secondly, we consider the case of central minimum, where $r_{-}=0$ and $n_{+}>0$ , corresponding to a flow that becomes singular at the axis of symmetry with the stretching rate exhibiting a local minimum at the minimum point $r_{0}=r_{-}$. These asymptotic behaviors are summarized in table 6. A critical threshold emerges at $n_{+}=4$, which we refer to as the blowup threshold. This means that the 3D Euler fluid under Lagrangian description remains regular if the local power-law approximation of the initial stretching rate satisfies $n_{+}\geq 4$, but develops a finite-time singularity when $n_{+}<4$.
@@
  • b.2)
    Central Minimum: Asymptotic Behavior of Plane Vorticity ($\omega$) and Vertical Pathline ($z$)
@@
    • The dependence on the cylinder radius is given by $\omega,z\propto R^{\frac{8}{4-n}-2}=R^{\frac{2n}{4-n}}$, showing that, for $n\in(0,4)$, larger radii slow down the decay of these quantities to zero.
@@
c) Non-central minimum. Thirdly, we consider the case of non-central minimum, where the minimum point of the initial stretching rate lies away from the axis of symmetry, i.e., $r_{-}\in(0,R]$, and $\gamma_{0}(r_{0})$ is continuous near this point with positive exponents $n_{\pm}\in\mathbb{R}_{+}$. The asymptotic behavior of the vorticity coordinate $(\omega,\gamma)$ and vertical pathline component $z$ in this scenario is summarized in table 7.
@@
These findings confirm that the local structure of the initial stretching rate $\gamma_{0}(r_{0})$ near its minimum plays a dominant role in determining whether and how the cylindrical Gibbon flow develops a finite-time singularity. The critical thresholds $n=1$ (blowup-rate threshold) and $n=2$ (blowup threshold) define distinct regimes of behavior:
@@
Summary of key features across central and non-central blowups. Now, we collect the following key observations regarding the relative strength of blowups, regardless of whether the blowup is central or non-central:
@@
  • For $n\in[2,4)$, blowup occurs only when the minimum point lies at the center; otherwise, for a non-central minimum, the solution remains regular;
@@
regardless of whether the blowup occurs at the center or away from it.
In addition to $r_{-}$, another parameter influencing the asymptotic behavior is $\delta$}, defined in equation (74), which quantifies the asymmetry between the left and right exponents $n_{-}$ and $n_{+}$ of $\gamma_{0}(r_{0})$ near its minimum. Since $\delta$} also appears in the denominator of $K_{1^{\pm}}$, it affects the blowup dynamics in the interval $n\in(1,2)$. Specifically, flows with asymmetric initial data—where $n_{-}\neq n_{+}$—exhibit faster blowup than those with symmetric profiles ($n_{-}=n_{+}$). In other words, a non-smooth or asymmetric profile of the initial stretching rate at the minimum point accelerates the formation of the singularity.
We emphasize a crucial physical distinction between central and non-central minimum scenarios. While the central minimum occurs on a point, the non-central minimum occurs on a ring, namely a degenerate case when compared to the point case. The ring case, due to the axisymmetry, has a perfectly “flat” profile in the azimuthal direction, which therefore changes the asymptotics in favor of more regular behavior, hence the required parameter $n$ for regular behavior is smaller compared with the central minimum case.
Finally we briefly address the question: What happens when the location of the minimum of the initial stretching rate is shared by several rings, or by a combination of rings and the central point? Namely, when the set $\{x\in\Omega\quad|\quad\gamma_{0}(x)=\min\gamma_{0}\}$ is the union of two or more elementary sets of the form $\{x\in\Omega\quad|\quad|x|=r_{-}\}$, with $r_{-}\geq 0$. In such a case, the dominant ring (or center point) is the one for which the integral $\Lambda_{1}$ grows faster asymptotically as $t\to T_{*}$. Thus, for example, if the minima of $\gamma_{0}$ were at a ring with $n=2$ and at the center with $n=2$, then the asymptotic behavior of the system would be dominated by the ring with $n=2$ because the integral $\Lambda_{1}$ would diverge due to the stretching rate’s profile at the ring, and thus the singularity time would still be $T_{*}=\infty$, even though a solitary initial minimum at the center with $n=2$ would lead to a finite-time singularity. As another example, a ring with $n=1$ combined with a center point with $n=4$ would lead to no finite-time singularity because the integral $\Lambda_{1}$ would diverge due to the stretching rate’s profile at the center, even though a solitary initial minimum at the ring with $n=1$ would lead to a finite-time singularity.